A clinical trial was conducted to test the effectiveness of a drug for treating insomnia in older subjects. Before​ treatment, 15 subjects had a mean wake time of 103.0 min. After​ treatment, the 15 subjects had a mean wake time of 97.5 min and a standard deviation of 21.9 min. Assume that the 15 sample values appear to be from a normally distributed population and construct a 95​% confidence interval estimate of the mean wake time for a population with drug treatments. What does the result suggest about the mean wake time of 103.0 min before the​ treatment? Does the drug appear to be​ effective?
Answer #1

Confidence Interval
X̅ ± t(α/2, n-1) S/√(n)
t(α/2, n-1) = t(0.05 /2, 15- 1 ) = 2.145
97.5 ± t(0.05/2, 15 -1) * 21.9/√(15)
Lower Limit = 97.5 - t(0.05/2, 15 -1) 21.9/√(15)
Lower Limit = 85.3722
Upper Limit = 97.5 + t(0.05/2, 15 -1) 21.9/√(15)
Upper Limit = 109.6278
95% Confidence interval is ( 85.3722 , 109.6278 )

Since the value of mean wake time of 103.0 min lies in the interval, hence we can conclude that drug does not appear to be effective.